# María Isabel Herrero1, Gabriela Jeronimo2,3,4 and Juan Sabia3,4

### v1

@@ 1 Introduction
When solving arbitrary polynomial systems, one may be interested in characterizing not only the isolated solutions but also the components of higher dimensions of the affine variety defined by the system. Early work in this direction was presented in [6] and [12], where the authors introduced new algorithms to compute the equidimensional decomposition of an affine variety. Later progress on this topic can be found in [10], [24], [28] and [22], with shorter running time algorithms relying on probabilistic methods. More recently, a new equidimensional decomposition algorithm whose efficiency has been shown in practice was presented in [9].
The equidimensional decomposition problem has also been studied in the sparse setting in [19], both from the theoretical and the algorihtmic points of view. For generic sparse systems with prescribed supports, the paper presents combinatorial conditions characterizing the equidimensional components of the algebraic variety defined by the system in the affine space, which are applied to design an efficient probabilistic algorithm to compute them. Also, some advances towards the algorithmic equidimensional decomposition for sparse systems with arbitrary coefficients were made in [19], but the problem was not completely solved.
In this paper, we design a symbolic probabilistic algorithm that characterizes completely the equidimensional components of the affine variety defined by an arbitrary sparse system. This characterization is given by means of witness sets of the equidimensional components of the variety, which is a well-known approach in numerical algebraic geometry (see, for instance, [34, Chapter 13]). For an equidimensional variety $W\subset\mathbb{A}^{n}$ of dimension $k$, a witness set of $W$ is a finite set of points characterizing $W$ obtained by intersecting $W$ with a generic linear variety of co-dimension $k$. Similarly to the notion of lifting fibers in the symbolic computation framework (see [28]), witness sets enables one to perform geometric computations related to the variety (see also [35]).
@@
Our approach is based on deformation techniques, which are widely used in numerical algebraic geometry for solving zero-dimensional polynomial systems (see, for instance, [34], [2]). Homotopic deformation methods have also been applied in symbolic algorithms. In [13], the authors introduced a new symbolic procedure for the computation of isolated points of algebraic varieties that significantly improved the previous complexity bounds. Combined with a symbolic Newton-Hensel lifting, these techniques lead to more efficient symbolic algorithms also solving some related elimination problems (see, for example, [14], [17], [15], [32]). Some algorithms for equidimensional decomposition of varieties also rely on these methods (see [24], [28], [22]). Algorithms designed specifically for the sparse polynomial setting are based on polyhedral deformations, which take into account the Newton polytopes of the input polynomials (see [36], [20], [21], [23], [18], [19]).
@@ 2.1 Sparse polynomial systems
Let $x=(x_{1},\dots,x_{n})$ be a family of indeterminates over a field $\mathbb{K}$ and $\mathbb{K}[x]=\mathbb{K}[x_{1},\dots,x_{n}]$ be the ring of polynomials in the variables $x$ with coefficients in $\mathbb{K}$. Throughout this paper, we will work with fields with characteristic 00, mainly with $\mathbb{K}=\mathbb{Q}$ or $\mathbb{K}=\mathbb{Q}(t)$ (the field of fractions of the polynomial ring in one variable with coefficients in $\mathbb{Q}$).
@@ 2.2 Affine varieties, equidimensional decomposition and witness sets
called a witness set of $W$. All witness sets of an equidimensional variety $W$ have the same cardinality, which is the degree of $W$, denoted by $\deg(W)$. Moreover, $\deg(W)$ is the maximum number of points in the intersection of $W$ with a linear affine variety of codimension $k$, provided that the intersection is finite (see [16]). The notion of witness set was introduced in the numerical algebraic geometry framework as a way to encode positive dimensional components of algebraic varieties. Indeed, if $W$ is an equidimensional component of the algebraic variety defined by a polynomial system ${\boldsymbol{f}}$, the data $(\mathcal{P},L,{\boldsymbol{f}})$ characterizes $W$ (see [34, Chapter 13]).
@@ 2.3 Complexity estimates
Starting from the foundational work of Giusti and Heintz (see [13]), straight-line programs showed to be effective in the construction of algorithms to solve many algebraic and geometric problems. This is the reason why we use this way of encoding in our intermediate computations.
@@
Multiplication and division with remainder of polynomials in $k[Y]$ of degrees bounded by $d$ can be computed with $O(M(d)/\log(d))$ arithmetic operations (see [37, Corollary 7.19 and Theorem 9.6]). The fast Extended Euclidean Algorithm (see [37, Section 11]) computes the gcd of two polynomials $f,g\in k[Y]$ of degrees bounded by $d$ within complexity $O(M(d))$. We compute the gcd of polynomials $h_{0},\dots,h_{m}\in k[Y]$ of degrees bounded by $d$ either in the standard recursive way, which takes $O(mM(d))$ operations, or in a probabilistic way, as $\gcd(h_{0},\sum_{1\leq j\leq m}c_{j}h_{j})$ for randomly chosen constants $c_{1},\dots,c_{m}$ (see [37, Algorithm 6.45]), within complexity $O(md+M(d))$ (see Section 4 for the notion of probabilistic algorithm we use).
Multipoint evaluation and interpolation of a polynomial $f\in k[Y]$ of degree $d$ at $d+1$ distinct points can be achieved within complexity $O(M(d))$ (see [37, Corollaries 10.8 and 10.12]). For a rational function $f=p/q$, where $p,q\in k[Y]$ have degrees bounded by $d$, we compute the dense representation of $p$ and $q$ from its Taylor series expansion up to order $2d$ using Padé approximation with $O(M(m))$ arithmetic operations in $k$ (see [37, Section 5.9]).
The sequence of subresultant polynomials of two polynomials $f,g\in k[Y]$ of degrees bounded by $d$ (in particular, their resultant) can be computed with $O(M(d))$ operations in $k$ (see [37, Corollary 11.19]).
In addition to manipulating polynomials, we will need to perform matrix computations. We denote $\Omega$ the exponent of the complexity of matrix multiplication, that is, a real number such that the product of two $n\times n$ matrices can be computed within complexity $O(n^{\Omega})$ (see [37, Section 12.1]). For practical computations, Strassen’s multiplication algorithm with $\Omega=\log 7$ can be used. The complexity of the computation of the determinant, the inverse and the characteristic polynomial of an $n\times n$ matrix can also be computed within $O(n^{\Omega})$ arithmetic operations.
@@ 3.1 A single point
Fix $k$, with $11\leq k\leq n-$ 1. Take $\ell_{1},\dots,\ell_{k}$ generic linear polynomials in $\mathbb{Q}[x_{1},\dots,x_{n}]$ and set
$L_{0}$ =$=\{{\boldsymbol{x}}\in\mathbb{A}^{n}\mid\ell_{i}({\boldsymbol{x}})-\ell_{i}({\boldsymbol{p}})=0\ \hbox{for all }1\leq i\leq k\}$
$L_{1}$ =$=\{{\boldsymbol{x}}\in\mathbb{A}^{n}\mid\ell_{i}(x)=0\ \hbox{for all }1\leq i\leq k\}$
@@
  • (H1)
    $V_{j}\cap L_{1}$ is either empty or $(j-k)$-equidimensional for all $j\geq k$, and $V_{j}\cap L_{1}=\emptyset$ for all $j<k$.
  • (H2)
@@
  • (H3)
    $V_{j}\cap L_{0}$ is either empty or $(j-k)$-equidimensional for all $j\geq k$, and $V_{j}\cap L_{0}$ is either empty or equal to $\{{\boldsymbol{p}}\}$ for all $j<k$.
@@
$\mathcal{L}$ =$=\{({\boldsymbol{x}},s)\in\mathbb{A}^{n+1}\mid\ell_{i}({\boldsymbol{x}})-(1-s)\ell_{i}({\boldsymbol{p}})=0\hbox{ for all }1\leq i\leq k\}$
@@
  • Proof.
    Let $V=\bigcup_{W}W$ be the irreducible decomposition of $V$. Then, $\widehat{V}\cap\mathcal{L}=\bigcup_{W}(\widehat{W}\cap\mathcal{L})$, where $\widehat{W}=W\times\mathbb{A}^{1}$ for each $W$. In particular, each irreducible component $C$ of $\mathcal{Z}$ is an irreducible component of $\widehat{W}\cap\mathcal{L}$ for some irreducible component $W$ of $V$.
@@
The previous lemmas enable us to detect the equidimensional component of $V$ of the largest dimension containing a point ${\boldsymbol{p}}\in V$. Consider $\ell_{1},\dots,\ell_{n-1}$ generic linear polynomials in $\mathbb{Q}[x_{1},\dots,x_{n}]$ so that assumptions (H1), (H2) and (H3) are met for every $11\leq k\leq n-$ 1.
For each $k$, with $11\leq k\leq n-$ 1, consider the variety $\mathcal{Z}^{(k)}\subset\mathbb{A}^{n+1}$ defined from $V$, $\ell_{1},\dots,\ell_{k}$ and ${\boldsymbol{p}}$ as above, and its equidimensional decomposition $\mathcal{Z}^{(k)}=\bigcup_{h=1}^{n-k}\mathcal{Z}^{(k)}_{h}$.
@@
  • Proof.
    By Lemma 1 for $k=n-1$, we have that $\mathcal{Z}^{(n-1)}=\mathcal{Z}_{1}^{(n-1)}\subset V_{n-1}\times\mathbb{A}^{1}$.
    Assume now that ${\boldsymbol{p}}\notin\bigcup_{j>k}V_{j}$ for a fixed $k$ (this holds, in particular, for $k=n-1$). Since ${\boldsymbol{p}}\in L_{0}$, Lemma 2 ensures that if $p\in V_{k}$, then $({\boldsymbol{p}},0)\in\mathcal{Z}_{1}^{(k)}\subset\mathcal{Z}^{(k)}$. On the other hand, due to Lemma 1, $\mathcal{Z}^{(k)}\subset\bigcup_{j\geq k}V_{j}\times\mathbb{A}^{1}$. Then, if $({\boldsymbol{p}},0)\in\mathcal{Z}^{(k)}$, it follows that ${\boldsymbol{p}}\in V_{k}$. $\square$
@@ 3.2 A finite set of points
For every ${\boldsymbol{p}}\in\mathcal{P}$, let $\theta_{{\boldsymbol{p}}}\in\mathbb{C}$ such that $\theta_{{\boldsymbol{p}}}\neq\theta_{{\boldsymbol{p}}^{\prime}}$ for ${\boldsymbol{p}}\neq{\boldsymbol{p}}^{\prime}$. As in Section 3.1, fix $k$ with $11\leq k\leq n-$ 1, take $\ell_{1},\dots,\ell_{k}$ (generic) linear affine functions satisfying (H1), (H2) and (H3), and consider
@@
We will now consider the previous construction for every $11\leq k\leq n-$ 1. Let $\ell_{1},\dots,\ell_{n-1}$ be generic linear polynomials in $\mathbb{Q}[x_{1},\dots,x_{n}]$ so that assumptions (H1), (H2), (H3) and (H4) are met for every $11\leq k\leq n-$ 1. For each $k$, let $\mathcal{Z}_{\mathcal{P}}^{(k)}\subset\mathbb{A}^{n+2}$ be a variety defined from the variety $V$, the finite set $\mathcal{P}$, and $\ell_{1},\dots,\ell_{k}$ as before.
@@ 4.1 Geometric resolutions
    The mixed volume in the right hand side of the above inequality counts the number of common zeros of a generic sparse system of polynomials in $n+1$ variables with supports $\{0,e_{1},\dots,e_{n},de_{0}\}$ and ($(\{0\}\times C_{i})\cup\{de_{0}\}$ for $i=1,\dots,n$; that is, a system of the form
@@ 4.2 Algorithmic deformation to classify points
Otherwise, by taking $n$ generic linear combinations of $f_{1},\dots,f_{m}$, we may obtain a polynomial system $\widetilde{{\boldsymbol{f}}}=(\widetilde{f}_{1},\dots,\widetilde{f}_{n})$ such that $V(\widetilde{{\boldsymbol{f}}})$ has the same equidimensional components as $V$ except, possibly, for the zero-dimensional one, which may contain some points not lying in $V$. Thus, we may apply the algorithm to $\widetilde{f}_{1},\dots,\widetilde{f}_{n}$ and remove the additional points where the system ${\boldsymbol{f}}$ does not vanish. We point out that the support of the polynomials $\widetilde{f}_{1},\dots,\widetilde{f}_{n}$ is the union of all the supports of $f_{1},\dots,f_{m}$ and so, mixed volumes associated to this set will appear in the complexities of our algorithms.
@@
For every ${\boldsymbol{p}}\in\mathcal{P}$, let $\theta_{{\boldsymbol{p}}}=\ell_{0}({\boldsymbol{p}})$. Then, $\theta_{\boldsymbol{p}}\neq\theta_{{\boldsymbol{p}}^{\prime}}$ for ${\boldsymbol{p}}\neq{\boldsymbol{p}}^{\prime}$ as required in Section 3.2. With this choice, for a fixed $k$ with $11\leq k\leq n-$ 1, we take linear affine polynomials $\ell_{1},\dots,\ell_{k}\in\mathbb{Q}[x_{1},\dots,x_{n}]$ randomly and consider the algebraic varieties defined as in Section 3.2:
@@
In the sequel, SparseSolving refers to the main algorithm from [23] adapted to deal with polynomials with coefficients in $\mathbb{Q}(s)$. It computes a finite set of points containing the isolated points in $\overline{\mathbb{Q}(s)}^{n}$ of the variety defined by $n$ input polynomials in $n$ variables. In order to keep a better control of the degrees in the parameter $s$, we modify slightly the last step of the algorithm so that the output is in Kronecker form. Intermediate computations are done with truncated power series in $s$ and rational functions are reconstructed by means of Padé approximation.
@@
INPUT: sparse polynomials $f_{1},\dots,f_{n}\in\mathbb{Q}[x_{1},\dots,x_{n}]$ defining a variety $V\subset\mathbb{A}^{n}$, univariate polynomials $M,v_{1},\dots,v_{n}$ in $\mathbb{Q}[x_{0}]$ representing a finite set of points $\mathcal{P}\subset V$ in shape-lemma form, an integer $k$ with $10\leq k\leq n-$ 1.
@@
    Apply the algorithm SparseSolving to the system ${\boldsymbol{F}}=(F_{0},F_{1},\dots,F_{n})$ to compute a geometric resolution $(m(Y),{\boldsymbol{w}}(Y))$ in $\mathbb{Q}(s)[Y]$ of a finite set of points of the variety $V({\boldsymbol{F}})\subset\mathbb{A}^{n+1}(\overline{\mathbb{Q}(s)})$ containing its isolated points.
@@
    Compute:
    • for $l=0,\dots,n+k$, $S_{l}^{H}(T,x_{0},x)=T^{\delta}S_{l}\left(\frac{1}{T}(x_{0},x)\right)$,
@@
      Compute $\gcd(m_{0},m_{0}^{\prime})$ and polynomials $P_{0}$, $P_{1}\in\mathbb{Q}[Y]$ such that $P_{0}m_{0}+P_{1}m_{0}^{\prime}=\gcd(m_{0},m_{0}^{\prime})$.
@@
Note that, since the polynomials ${\boldsymbol{S}}$ define the variety $\widehat{V}_{\mathcal{P}}\cap\mathcal{L}_{\mathcal{P}}$ (see (2)) and $\mathcal{Z}_{\mathcal{P}}$ is the union of all irreducible components $\mathcal{C}$ of this variety with $I(\mathcal{C})\cap\mathbb{Q}[s]=\{0\}$, the equidimensional components of $V({\boldsymbol{S}})\subset\mathbb{A}^{n+1}(\overline{\mathbb{Q}(s)})$ are in one to one correspondence with the equidimensional components of $\mathcal{Z}_{\mathcal{P}}\subset\mathbb{A}^{n+2}$. In particular, the set of isolated points of $V({\boldsymbol{S}})$ corresponds to the equidimensional component $\mathcal{Z}_{\mathcal{P},1}$ of dimension $1$. Now, since the finite set $\overline{\mathcal{Q}}$ contains the isolated points of $V({\boldsymbol{S}})$, it corresponds to an equidimensional subvariety $\mathcal{W}$ of $\mathcal{Z}_{\mathcal{P}}$ of dimension $1$ that contains $\mathcal{Z}_{\mathcal{P},1}$. Step 6 of the algorithm produces univariate polynomials associated with the intersection $\mathcal{W}\cap\{s=0\}$ provided that the linear form used in the computations of Step 4 is well-separating for $\overline{\mathcal{Q}}$. More precisely, following [1, Section 12.5], the algorithm determines the limits when $s\to 0$ of the bounded points of $\overline{\mathcal{Q}}$.
@@
$\mathcal{B}_{0}$ =($=(\{0\}\times\Delta_{n})\cup\{(\alpha,\mathbf{0}_{n}):0\leq\alpha\leq\delta\}\subset(\mathbb{Z}_{\geq 0})^{n+1},$
$\mathcal{B}_{l}$ =($=(\{0\}\times(\mathcal{A}_{l}\cup\Delta_{n}))\cup\{(\alpha,\mathbf{0}_{n}):0\leq\alpha\leq\delta-1\}\subset(\mathbb{Z}_{\geq 0})^{n+1},\quad\hbox{for }l=1,\dots,n.$
@@
Algorithm SparseSolving works in two steps: first, it computes a geometric resolution of a generic sparse system of polynomials with supports $\mathcal{B}_{0},\dots,\mathcal{B}_{n}$ and then, it makes a homotopic deformation to obtain a geometric resolution of the isolated zeros of the given system $F_{0},\dots,F_{n}$. Up to a factor depending on the computation of a suitable subdivision of the family of supports, by [23, Proposition 5.13], the complexity of the first step (over $\mathbb{Q}$) is of order

### v2

@@ 1 Introduction
When solving arbitrary polynomial systems, one may be interested in characterizing not only the isolated solutions but also the components of higher dimensions of the affine variety defined by the system (for the best known complexity bounds for solving zero-dimensional systems see [37]). Early work in this direction was presented in [6] and [12], where the authors introduced new algorithms to compute the equidimensional decomposition of an affine variety. Later progress on this topic can be found in [10], [28], [24], [29] and [22], with shorter running time algorithms relying on probabilistic methods. More recently, a new equidimensional decomposition algorithm whose efficiency has been shown in practice was presented in [9].
The equidimensional decomposition problem has also been studied in the sparse setting in [19], both from the theoretical and the algorithmic points of view. For generic sparse systems with prescribed supports, the paper presents combinatorial conditions characterizing the equidimensional components of the algebraic variety defined by the system in the affine space, which are applied to design an efficient probabilistic algorithm to compute them. Also, some advances towards the algorithmic equidimensional decomposition for sparse systems with arbitrary coefficients were made in [19], but the problem was not completely solved, since the proposed algorithm provides partial information on the equidimensional components of the variety but it does not give a full description of them.
In this paper, we design a symbolic probabilistic algorithm that characterizes completely the equidimensional components of the affine variety defined by an arbitrary sparse system. This characterization is given by means of witness sets of the equidimensional components of the variety, which is a well-known approach in numerical algebraic geometry (see, for instance, [35, Chapter 13]). For an equidimensional variety $W\subset\mathbb{A}^{n}$ of dimension $k$, a witness set of $W$ is a finite set of points characterizing $W$ obtained by intersecting $W$ with a generic linear variety of co-dimension $k$. Similarly to the notion of lifting fibers in the symbolic computation framework (see [29]), witness sets enable one to perform geometric computations related to the variety (see also [36]).
@@
Our approach is based on deformation techniques, which are widely used in numerical algebraic geometry for solving zero-dimensional polynomial systems (see, for instance, [35], [2]). Homotopic deformation methods have also been applied in symbolic algorithms. In [13], the authors introduced a new symbolic procedure for the computation of isolated points of algebraic varieties that significantly improved the previous complexity bounds. Combined with a symbolic Newton-Hensel lifting, these techniques lead to more efficient symbolic algorithms also solving some related elimination problems (see, for example, [14], [17], [15], [33]). Some algorithms for equidimensional decomposition of varieties also rely on these methods (see [28], [24], [29], [22]). Algorithms designed specifically for the sparse polynomial setting are based on polyhedral deformations, which take into account the Newton polytopes of the input polynomials (see [38], [20], [21], [23], [18], [19]).
@@ 2.1 Sparse polynomial systems
Let $x=(x_{1},\dots,x_{n})$ be a family of indeterminates over a field $\mathbb{K}$ and $\mathbb{K}[x]=\mathbb{K}[x_{1},\dots,x_{n}]$ be the ring of polynomials in the variables $x$ with coefficients in $\mathbb{K}$. Throughout this paper, we will work with fields with characteristic 0, mainly with $\mathbb{K}=\mathbb{Q}$ or $\mathbb{K}=\mathbb{Q}(t)$ (the field of fractions of the polynomial ring in one variable with coefficients in $\mathbb{Q}$).
@@ 2.2 Affine varieties, equidimensional decomposition and witness sets
called a witness set of $W$. All witness sets of an equidimensional variety $W$ have the same cardinality, which is the degree of $W$, denoted by $\deg(W)$. Moreover, $\deg(W)$ is the maximum number of points in the intersection of $W$ with a linear affine variety of codimension $k$, provided that the intersection is finite (see [16]). The notion of witness set was introduced in the numerical algebraic geometry framework as a way to encode positive dimensional components of algebraic varieties. Indeed, if $W$ is an equidimensional component of the algebraic variety defined by a polynomial system $\boldsymbol{f}$}, the data $(\mathcal{P},L,{\boldsymbol{f}})$ characterizes $W$ (see [35, Chapter 13]).
@@ 2.3 Complexity estimates
Starting from the foundational work of Giusti and Heintz (see [13]), straight-line programs proved to be effective in the construction of algorithms to solve many algebraic and geometric problems. This is the reason why we use this way of encoding in our intermediate computations.
@@
Multiplication and division with remainder of polynomials in $k[Y]$ of degrees bounded by $d$ can be computed with $O(M(d)/\log(d))$ arithmetic operations (see [39, Corollary 7.19 and Theorem 9.6]). The fast Extended Euclidean Algorithm (see [39, Section 11]) computes the gcd of two polynomials $f,g\in k[Y]$ of degrees bounded by $d$ within complexity $O(M(d))$. We compute the gcd of polynomials $h_{0},\dots,h_{m}\in k[Y]$ of degrees bounded by $d$ either in the standard recursive way, which takes $O(mM(d))$ operations, or in a probabilistic way, as $\gcd(h_{0},\sum_{1\leq j\leq m}c_{j}h_{j})$ for randomly chosen constants $c_{1},\dots,c_{m}$ (see [39, Algorithm 6.45]), within complexity $O(md+M(d))$ (see Section 4 for the notion of probabilistic algorithm we use).
Multipoint evaluation and interpolation of a polynomial $f\in k[Y]$ of degree $d$ at $d+1$ distinct points can be achieved within complexity $O(M(d))$ (see [39, Corollaries 10.8 and 10.12]). For a rational function $f=p/q$, where $p,q\in k[Y]$ have degrees bounded by $d$, we compute the dense representation of $p$ and $q$ from its Taylor series expansion up to order $2d$ using Padé approximation with $O(M(m))$ arithmetic operations in $k$ (see [39, Section 5.9]).
Any subresultant polynomial of two polynomials $f,g\in k[Y]$ of degrees bounded by $d$ (in particular, their resultant) can be computed with $O(M(d))$ operations in $k$ (see [39, Corollary 11.19]).
In addition to manipulating polynomials, we will need to perform matrix computations. We denote $\Omega$ the exponent of the complexity of matrix multiplication, that is, a real number such that the product of two $n\times n$ matrices can be computed within complexity $O(n^{\Omega})$ (see [39, Section 12.1]). For practical computations, Strassen’s multiplication algorithm with $\Omega=\log 7$ can be used. The complexity of the computation of the determinant, the inverse and the characteristic polynomial of an $n\times n$ matrix can also be computed within $O(n^{\Omega})$ arithmetic operations.
@@ 3.1 A single point
Fix $k$, with $1\leq k\leq n-1$. Take $\ell_{1},\dots,\ell_{k}$ generic linear polynomials in $\mathbb{Q}[x_{1},\dots,x_{n}]$ and set
$L_{0}$ $=\{{\boldsymbol{x}}\in\mathbb{A}^{n}\mid\ell_{i}({\boldsymbol{x}})-\ell_{i}({\boldsymbol{p}})=0\ \hbox{for all }1\leq i\leq k\}$
$L_{1}$ $=\{{\boldsymbol{x}}\in\mathbb{A}^{n}\mid\ell_{i}(x)=0\ \hbox{for all }1\leq i\leq k\}$
@@
$V_{j}\cap L_{1}$ is either empty or (jk)(j-k)-equidimensional for all $j\geq k$, and $V_{j}\cap L_{1}=\emptyset$ for all $j<k$.
@@
$V_{j}\cap L_{0}$ is either empty or (jk)(j-k)-equidimensional for all $j\geq k$, and $V_{j}\cap L_{0}$ is either empty or equal to $\{{\boldsymbol{p}}\}$ for all $j<k$.
@@
$\mathcal{L}$ $=\{({\boldsymbol{x}},s)\in\mathbb{A}^{n+1}\mid\ell_{i}({\boldsymbol{x}})-(1-s)\ell_{i}({\boldsymbol{p}})=0\hbox{ for all }1\leq i\leq k\}$
@@
[…]
@@
The previous lemmas enable us to detect the equidimensional component of $V$ of the largest dimension containing a point ${\boldsymbol{p}}\in V$. Consider $\ell_{1},\dots,\ell_{n-1}$ generic linear polynomials in $\mathbb{Q}[x_{1},\dots,x_{n}]$ so that assumptions (H1), (H2) and (H3) are met for every $1\leq k\leq n-1$.
For each $k$, with $1\leq k\leq n-1$, consider the variety $\mathcal{Z}^{(k)}\subset\mathbb{A}^{n+1}$ defined from $V$, $\ell_{1},\dots,\ell_{k}$ and $\boldsymbol{p}$} as above, and its equidimensional decomposition $\mathcal{Z}^{(k)}=\bigcup_{h=1}^{n-k}\mathcal{Z}^{(k)}_{h}$.
@@
By Lemma 1 for k=n1k=n-1, we have that $\mathcal{Z}^{(n-1)}=\mathcal{Z}_{1}^{(n-1)}\subset V_{n-1}\times\mathbb{A}^{1}$.
Assume now that ${\boldsymbol{p}}\notin\bigcup_{j>k}V_{j}$ for a fixed $k$ (this holds, in particular, for k=n1k=n-1). Since ${\boldsymbol{p}}\in L_{0}$, Lemma 2 ensures that if $p\in V_{k}$, then $({\boldsymbol{p}},0)\in\mathcal{Z}_{1}^{(k)}\subset\mathcal{Z}^{(k)}$. On the other hand, due to Lemma 1, $\mathcal{Z}^{(k)}\subset\bigcup_{j\geq k}V_{j}\times\mathbb{A}^{1}$. Then, if $({\boldsymbol{p}},0)\in\mathcal{Z}^{(k)}$, it follows that ${\boldsymbol{p}}\in V_{k}$. $\square$
@@ 3.2 A finite set of points
For every ${\boldsymbol{p}}\in\mathcal{P}$, let $\theta_{{\boldsymbol{p}}}\in\mathbb{C}$ such that $\theta_{{\boldsymbol{p}}}\neq\theta_{{\boldsymbol{p}}^{\prime}}$ for ${\boldsymbol{p}}\neq{\boldsymbol{p}}^{\prime}$. As in Section 3.1, fix $k$ with $1\leq k\leq n-1$, take $\ell_{1},\dots,\ell_{k}$ (generic) linear affine functions satisfying (H1), (H2) and (H3), and consider
@@
We will now consider the previous construction for every $1\leq k\leq n-1$. Let $\ell_{1},\dots,\ell_{n-1}$ be generic linear polynomials in $\mathbb{Q}[x_{1},\dots,x_{n}]$ so that assumptions (H1), (H2), (H3) and (H4) are met for every $1\leq k\leq n-1$. For each $k$, let $\mathcal{Z}_{\mathcal{P}}^{(k)}\subset\mathbb{A}^{n+2}$ be a variety defined from the variety $V$, the finite set $\mathcal{P}$, and $\ell_{1},\dots,\ell_{k}$ as before.
@@ 4.1 Geometric resolutions
The mixed volume in the right hand side of the above inequality counts the number of common zeros of a generic sparse system of polynomials in $n+1$ variables with supports $\{0,e_{1},\dots,e_{n},de_{0}\}$ and $(\{0\}\times C_{i})\cup\{de_{0}\}$ for $i=1,\dots,n$; that is, a system of the form
@@ 4.2 Algorithmic deformation to classify points
[…]
@@
For every ${\boldsymbol{p}}\in\mathcal{P}$, let $\theta_{{\boldsymbol{p}}}=\ell_{0}({\boldsymbol{p}})$. Then, $\theta_{\boldsymbol{p}}\neq\theta_{{\boldsymbol{p}}^{\prime}}$ for ${\boldsymbol{p}}\neq{\boldsymbol{p}}^{\prime}$ as required in Section 3.2. With this choice, for a fixed $k$ with $1\leq k\leq n-1$, we take linear affine polynomials $\ell_{1},\dots,\ell_{k}\in\mathbb{Q}[x_{1},\dots,x_{n}]$ randomly and consider the algebraic varieties defined as in Section 3.2:
@@
In the sequel, SparseSolving refers to the main algorithm from [23] adapted to deal with polynomials with coefficients in $\mathbb{Q}(s)$. Given an input system of $n$ polynomials in $n$ variables in sparse representation, it computes a finite set of points containing the isolated points in $\overline{\mathbb{Q}(s)}^{n}$ of the variety defined by the system. In order to keep a better control of the degrees in the parameter $s$, we modify slightly the last step of the algorithm so that the output is in Kronecker form. Intermediate computations are done with truncated power series in $s$ and rational functions are reconstructed by means of Padé approximation.
@@
Following the notation in [1, Section 12.5], for a polynomial $p\in\mathbb{Q}(s)[Y]$, we write $\lim\limits_{s\to 0}(p)$ for the polynomial which is obtained from $p(Y)$ by evaluating at $s=0$ the coefficients of the polynomial $s^{-o}p(Y)$, where $o$ is the minimum of the orders of the coefficients of $p(Y)$.
@@
INPUT: polynomials $f_{1},\dots,f_{n}\in\mathbb{Q}[x_{1},\dots,x_{n}]$ in sparse representation defining a variety $V\subset\mathbb{A}^{n}$, univariate polynomials $M,v_{1},\dots,v_{n}$ in $\mathbb{Q}[x_{0}]$ representing a finite set of points $\mathcal{P}\subset V$ in shape-lemma form, an integer $k$ with $0\leq k\leq n-1$.
@@
    Apply the algorithm SparseSolving to the system ${\boldsymbol{F}}=(F_{0},F_{1},\dots,F_{n})$ to compute a geometric resolution $(m(Y),{\boldsymbol{w}}(Y))$ in $\mathbb{Q}(s)[Y]$ of a finite set of points $\widetilde{\mathcal{Q}}$ of the variety $V({\boldsymbol{F}})\subset\mathbb{A}^{n+1}(\overline{\mathbb{Q}(s)})$ containing its isolated points.
@@
    Compute a geometric resolution of $\widetilde{\mathcal{Q}}\cap V(S_{0},\dots,S_{n+k})$:
    • $\delta=\max\{\deg(f_{1}),\dots,\deg(f_{n}),\deg(M)\}$,
    • for $l=0,\dots,n+k$, $S_{l}^{H}(T,x_{0},x)=T^{\delta}S_{l}\left(\frac{1}{T}(x_{0},x)\right)$ (here $T$ is a new variable),
@@
      Compute $\gcd(m_{0}(Y),m_{0}^{\prime}(Y))$ and polynomials $P_{0}(Y)$, $P_{1}(Y)\in\mathbb{Q}[Y]$ such that $P_{0}(Y)m_{0}(Y)+P_{1}(Y)m_{0}^{\prime}(Y)=\gcd(m_{0}(Y),m_{0}^{\prime}(Y))$.
@@
Note that, since the polynomials $\boldsymbol{S}$} define the variety $\widehat{V}_{\mathcal{P}}\cap\mathcal{L}_{\mathcal{P}}$ (see (2)) and $\mathcal{Z}_{\mathcal{P}}$ is the union of all irreducible components $\mathcal{C}$ of this variety with $I(\mathcal{C})\cap\mathbb{Q}[s]=\{0\}$, the equidimensional components of $V({\boldsymbol{S}})\subset\mathbb{A}^{n+1}(\overline{\mathbb{Q}(s)})$ are in one to one correspondence with the equidimensional components of $\mathcal{Z}_{\mathcal{P}}\subset\mathbb{A}^{n+2}$. In particular, the set of isolated points of $V({\boldsymbol{S}})$ corresponds to the equidimensional component $\mathcal{Z}_{\mathcal{P},1}$ of dimension $1$. Now, since the finite set $\overline{\mathcal{Q}}$ contains the isolated points of $V({\boldsymbol{S}})$, it corresponds to an equidimensional subvariety $\mathcal{W}$ of $\mathcal{Z}_{\mathcal{P}}$ of dimension $1$ that contains $\mathcal{Z}_{\mathcal{P},1}$.
Step 6 of the algorithm produces univariate polynomials associated with the intersection $\mathcal{W}\cap\{s=0\}$ provided that the linear form used in the computations of Step 4 is well-separating for $\overline{\mathcal{Q}}$. More precisely, following [1, Section 12.5], the algorithm determines the limits when $s\to 0$ of the bounded points of $\overline{\mathcal{Q}}$.
@@
$\mathcal{B}_{0}$ $=(\{0\}\times\Delta_{n})\cup\{(\alpha,\mathbf{0}_{n}):0\leq\alpha\leq\delta\}\subset(\mathbb{Z}_{\geq 0})^{n+1},$
$\mathcal{B}_{l}$ $=(\{0\}\times(\mathcal{A}_{l}\cup\Delta_{n}))\cup\{(\alpha,\mathbf{0}_{n}):0\leq\alpha\leq\delta-1\}\subset(\mathbb{Z}_{\geq 0})^{n+1},\quad\hbox{for }l=1,\dots,n.$
@@
Algorithm SparseSolving works in two stages: first, it computes a geometric resolution of a generic sparse system of polynomials with supports $\mathcal{B}_{0},\dots,\mathcal{B}_{n}$ and then, it makes a homotopic deformation to obtain a geometric resolution of the isolated zeros of the given system $F_{0},\dots,F_{n}$. Up to a factor depending on the computation of a suitable subdivision of the family of supports, by [23, Proposition 5.13], the complexity of the first stage (over $\mathbb{Q}$) is of order